OpenStudy (anonymous):

What is the 15th term in the sequence? 3 , 12 , 48 , 192 , 768, . . .

OpenStudy (anonymous):

Do you know the first step to solving this problem?

OpenStudy (anonymous):

First you apply the formula, do you know what that is?

OpenStudy (anonymous):

an+a1+(n-1)d

OpenStudy (anonymous):

this is a geometric sequence

OpenStudy (anonymous):

no arithmetic

hartnn (hartnn):

how do you figure out whether the series is arithmetic or geometric?

OpenStudy (anonymous):

this is a geometric sequence with a common ration of 4

hartnn (hartnn):

@Morameme1 do you know how?

OpenStudy (anonymous):

An arithmetic sequence goes from one term to the next by always adding (or subtracting) the same value.

OpenStudy (anonymous):

A geometric sequence goes from one term to the next by always multiplying (or dividing) by the same value.

OpenStudy (anonymous):

why? because each term after the first is being multiply by the first number and a fix number?

OpenStudy (anonymous):

previous number by a fixed value

OpenStudy (anonymous):

You need to multiply by 4.

OpenStudy (anonymous):

Until you get to the 12th number, and add.

OpenStudy (anonymous):

therefore this will make it a geometric and not arithmetic :D

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

Add them all together. Like: 3*4= 12; 12*4= 48; 48*4=192 and so on.. When you get to your 12th number, add them all up then you should get - 16777215

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

yea thx

OpenStudy (anonymous):

You're welcome.

OpenStudy (dayakar):

t1= 3 t2=12 t3= 48 t4= 192 observe t2/t1= t3/t2= t4/t3= 4 therefore common ratio (r) =4 the sequence is in geometic progression

OpenStudy (dayakar):

t15 =a*r^14 here a= first term = 3 r= 4 plug in the values and get t15